OpenStudy (anonymous):

Solving Rational Equations What is the product in simplest form? State any restrictions on the variable. 4a^5/7b^4 * 2b^2/2a^4

OpenStudy (anonymous):

\[\frac{ 4a ^{5} }{7b ^{4}}*\frac{ 2b ^{2} }{ 2a ^{4} }=\frac{ 4a ^{5}*a ^{-4} }{7b ^{4}*b ^{-2} }=\frac{ 4a ^{5-4} } {7b ^{4-2} }\] simplify it. Restriction is \[a \neq 0,b \neq0\]

OpenStudy (anonymous):

Thanks!